Using the definition of absolute value, evaluate "\begin{vmatrix}
 X-2/3 =3/4
 
\end{vmatrix}\mid"

1
Expert's answer
2021-11-30T11:54:45-0500

"|x-\frac{2}{3}|=\frac{3}{4}\\
\text{By using definition of absolute value, we have}\\
\implies x-\frac{2}{3}=\frac{3}{4} \space or\space
 -(x-\frac{2}{3})=\frac{3}{4} \\

\implies x=\frac{2}{3}+\frac{3}{4} \space or \space x= -\frac{3}{4}+\frac{2}{3}\\

\implies x=\frac{17}{12} \space or \space x= -\frac{1}{12}\\"